OpenStudy (anonymous):

Can someone help me finish separating this differential equation? I have y'=(sin5x)/(siny) dy/dx=(sin5x)/(siny) sinydy=sin5xdx integrate both sides -cosy=(1/2)x^2sin5 how do I get to the final explicit form?

OpenStudy (anonymous):

I might be missing something, but isn't \[\int\limits_{}^{} \sin(5x) = -\cos (5x) / 5\]

OpenStudy (anonymous):

regardless... you (I believe) to get it into explicit form, solve for y. Which means taking the arc cosine.

OpenStudy (anonymous):

okay, so my last question then, is where does C go? like, as in: arccos(1/5cos5x)+C or arccos(1/5cos5x+C)